OpenStudy (tennistar):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3.

OpenStudy (welshfella):

let f(x) = y y = 4(x + 3) - 5 y = 4x + 12 - 5 y = 4x + 7 now can you solve this for x?

OpenStudy (tennistar):

I just put the value of x in place of the x, correct?

OpenStudy (tennistar):

So that would be 4(3) + 7 = 19, yes?

OpenStudy (tennistar):

But I am supposed to find the inverse function @welshfella